# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix D Minimality and Attainability of HUC
and similarly the joint distribution $Q^{\mathtt{Slice}}$ has
[…]
The Rényi divergence of order $\alpha>1$ between these joint probability measures is
$\mathtt{D}_{\alpha}(\mathrm{P}^{\mathtt{Slice}}\|\mathrm{Q}^{\mathtt{Slice}})$ $=\frac{1}{\alpha-1}\log\sum_{\ell=1}^{m}\int f(\ell,\hat{z})^{\alpha}g(\ell,z)^{1-\alpha}dz$
$=\frac{1}{\alpha-1}\log\sum_{\ell=1}^{m}\omega_{\ell}\int p_{\ell}(z)^{\alpha}q_{\ell}(z)^{1-\alpha}dz$
$=\frac{1}{\alpha-1}\log\sum_{\ell=1}^{m}\omega_{\ell}\exp\bigl((\alpha-1)D_{\ell}\bigr)$
$=\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}),$
since
$D_{\ell}=\mathtt{D}_{\alpha}(u_{\ell}{}_{\#}P\|u_{\ell}{}_{\#}Q)=\frac{1}{\alpha-1}\log\int p_{\ell}(z)^{\alpha}q_{\ell}(z)^{1-\alpha}dz.$
Thus, we have
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})=\mathtt{D}_{\alpha}(\mathrm{P}^{\mathtt{Slice}}\|\mathrm{Q}^{\mathtt{Slice}}).$
[…]
$\mathtt{D}_{\alpha}(\mathrm{P}K\|\mathrm{Q}K)\leq\mathtt{D}_{\alpha}(\mathrm{P}\|\mathrm{Q}).$
When the kernel $\mathsf{K}$ is the sliced channel $\mathsf{Slice}$, we obtain
$\mathtt{D}_{\alpha}(\mathrm{P}^{\mathtt{Slice}}\|\mathrm{Q}^{\mathtt{Slice}})\leq\mathtt{D}_{\alpha}(\mathrm{P}\|\mathrm{Q}).$
That is,
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})\leq\mathtt{D}_{\alpha}(\mathrm{P}\|\mathrm{Q}).$
By combining Part 1 and Part 2, we can claim
@@
$\mathtt{ASD}^{\omega}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big)\leq\mathtt{JSD}^{\omega}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big)\leq\mathtt{D}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big),$
Appendix F Proof of Lemma 4.1
Fix a prior $\theta\in\Theta$, a secret pair $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$, a slice direction $u\in\mathcal{U}$, and an order $\alpha>1$. Define the 1-D projected pre-noise probability measures as
$P_{u}:=Psi^{u}_{\#}F^{\theta}_{s_{i}}\quad\textup{and}\quad Q_{u}:=\Psi^{u}_{\#}F^{\theta}_{s_{j}}.$
By definition of the per-slice $\infty$-Wasserstein distance,
$z_{\infty}^{u}:=W_{\infty}(P_{u},Q_{u})=\inf_{\pi\in\Pi(P_{u},Q_{u})}\operatorname*{ess\,sup}_{(z,z^{\prime})\sim\pi}|z-z^{\prime}|.$
Therefore, for every $\varepsilon>0$ there exists a coupling $\pi_{\varepsilon}\in\Pi(P_{u},Q_{u})$ such that
$|Z-Z^{\prime}|\leq z_{\infty}^{u}+\varepsilon\quad\text{for $\pi_{\varepsilon}$-almost every $(Z,Z^{\prime})$.}$ (28)
Step 1. We first represent the sliced mechanism as a mixture of shifts. Let $\widetilde{N}:=\Psi^{u}(N)$ be the projected noise with distribution $\Psi^{u}_{\#}\zeta$. With abuse of notation, we denote this 1-D noise distribution again by $\zeta$.
Conditioned on a pre-noise scalar $z\in\mathbb{R}$, the sliced mechanism output is $z+\widetilde{N}$, whose probability measure is the shifted noise distribution:
$\zeta_{z}(A):=\Pr(z+\widetilde{N}\in A)=\zeta(A-z),\quad\textup{for}\quad A\subseteq\mathbb{R}.$
Under secret $s_{i}$, the sliced output distribution is therefore
$\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}=\int\zeta_{z}P_{u}(dz).$
Similarly, under $s_{j}$,
$\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}=\int\zeta_{z^{\prime}}Q_{u}(dz^{\prime}).$
Using the coupling $\pi_{\varepsilon}$ between $P_{u}$ and $Q_{u}$, we can represent $\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}$ and $\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}$ as mixtures over pairs $(z,z^{\prime})$:
$\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}=\int\zeta_{z}\pi_{\varepsilon}(dz,dz^{\prime})\quad\textup{and}\quad\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}=\int\zeta_{z^{\prime}}\pi_{\varepsilon}(dz,dz^{\prime}).$
Step 2. Next, we obtain a pointwise Rényi divergence between shifted noises. Let $\zeta_{z}$ denote the distribution of $\widetilde{N}+z$ for $z\in\mathbb{R}$. Fix a pair $(z,z^{\prime})$ with $z\in\mathbb{R}$ and $z^{\prime}\in\mathbb{R}$. Since Rényi divergence is invariant under common bijective transformations, the corresponding shifted noise probability measures $\zeta_{z}$ and $\zeta_{z^{\prime}}$ satisfy
@@
$\mathtt{D}_{\alpha}(\zeta_{z}\|\zeta_{z^{\prime}})=\mathtt{D}_{\alpha}(\zeta_{z-z^{\prime}}\|\zeta).$
By the definition of the shift-Rényi envelope
$R_{\alpha}(\zeta,r):=\sup_{|a|\leq r}\mathtt{D}_{\alpha}(\zeta_{a}\|\zeta),$
we have, for every $z\in\mathbb{R}$ and $z^{\prime}\in\mathbb{R}$,
$\mathtt{D}_{\alpha}(\zeta_{z}\|\zeta_{z^{\prime}})\leq R_{\alpha}(\zeta,|z,z^{\prime}|).$
Combining with the coupling bound (28), we obtain the uniform pointwise bound
$\mathtt{D}_{\alpha}(\zeta_{z}\|\zeta_{z^{\prime}})\leq R_{\alpha}\bigl(\zeta,z_{\infty}^{u}\bigr),$ (29)
for $\pi_{\varepsilon}$-a.e. $(z,z^{\prime})$.
Step 3. We proceed with the full mixtures $\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}$ and $\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}$. Rényi divergence of order $\alpha>1$ is jointly quasi-convex in its arguments. That is, for any probability measure $\mu$ over an index set and any family $\{P_{\beta},Q_{\beta}\}$,
$\mathtt{D}_{\alpha}\left(\int P_{\beta}\mu(d\beta)\Big\|\int Q_{\beta}\mu(d\beta)\right)\leq\operatorname*{ess\,sup}_{\beta}\mathtt{D}_{\alpha}(P_{\beta}\|Q_{\beta}).$
For finite mixtures, this follows from the standard inequality
$\mathtt{D}_{\alpha}\left(\sum_{k}\lambda_{k}P_{k}\Big\|\sum_{k}\lambda_{k}Q_{k}\right)\leq\max_{k}\mathtt{D}_{\alpha}(P_{k}\|Q_{k}),$
and the general case can be obtained by approximating $\mu$ by finite-support measures (a standard argument for $f$-divergences).
In our setting we take $\beta=(z,z^{\prime})$, $\mu=\pi$ (the coupling from Step 1), and $P_{\beta}=\zeta_{z}$ and $Q_{\beta}=\zeta_{z^{\prime}}$. Then
$\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}=\int P_{\beta}\pi(dz,dz^{\prime}),\qquad\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}=\int Q_{\beta}\pi(dz,dz^{\prime}).$
Hence
$\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}\big\|\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}\bigr)\leq\operatorname*{ess\,sup}_{(z,z^{\prime})\sim\pi}\mathtt{D}_{\alpha}(\zeta_{z}\|\zeta_{z^{\prime}}).$
Combining this with the pointwise bound from Step 2,
$\mathtt{D}_{\alpha}(\zeta_{z}\|\zeta_{z^{\prime}})\leq R_{\alpha}(\zeta,z_{\infty}^{u})$
for $\pi$-a.e. $(z,z^{\prime})$, yields
@@
$\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}\big\|\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}\bigr)\leq R_{\alpha}(\zeta,z_{\infty}^{u}),$
which is the desired per-slice envelope bound. ∎
Appendix G Proof of Theorem 4.7
Fix $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$, and fix a slice profile $\{\mathcal{U},\omega\}$. Let $F^{\theta}_{s}$ be the distribution of the pre-noise output $f(X)$ given $(S=s,\theta)$, and let
$\mathrm{P}:=\mathcal{M}^{\theta}_{s_{i}}=F^{\theta}_{s_{i}}*\zeta\quad\textup{and}\quad\mathrm{Q}:=\mathcal{M}^{\theta}_{s_{j}}=F^{\theta}_{s_{j}}*\zeta.$
Step 1. We start by establishing a per-slice envelop bound using Lemma 4.1. For any direction $u\in\mathcal{U}$, applying Lemma 4.1 to the 1-D projection $\Psi^{u}(y)=\langle y,u\rangle$ gives, for all $\alpha>1$,
@@
$\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\ \big\|\ \Psi^{u}_{\#}\mathrm{Q}\bigr)\leq R_{\alpha}\bigl(\zeta,z^{u}_{\infty}(\theta,s_{i},s_{j})\bigr),$
where $z^{u}_{\infty}(\theta,s_{i},s_{j})$ is any upper bound on the 1-D shift between the pre-noise probability measures $\Psi^{u}_{\#}F^{\theta}_{s_{i}}$ and $\Psi^{u}_{\#}F^{\theta}_{s_{j}}$. By definition,
$z^{u}_{\infty}(\theta,s_{i},s_{j})\leq\tilde{\Delta}^{u}_{\infty}(\theta,s_{i},s_{j}):=W_{\infty}\bigl(\Psi^{u}_{\#}F^{\theta}_{s_{i}}\;\big\|\;\Psi^{u}_{\#}F^{\theta}_{s_{j}}\bigr),$
and $R_{\alpha}(\zeta,\cdot)$ is non-decreasing in its second argument. Thus, we have
@@
$\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}Q\bigr)\leq R_{\alpha}\bigl(\zeta,\tilde{\Delta}_{\infty}^{u}(\theta,s_{i},s_{j})\bigr)$ (30)
for all $u\in\mathcal{U}$.
@@
Step 2. Next, we consider Gaussian noise $\zeta=\mathcal{N}(0,\sigma^{2}I_{d})$. For any slice $u\in\mathcal{U}$, the projected noise $\Psi^{u}(N)$ is one-dimensional Gaussian $\mathcal{N}(0,\sigma^{2})$. For any shift $a\in\mathbb{R}$, the shifted probability measure is $\mathcal{N}(a,\sigma^{2})$, and the Rényi divergence of order $\alpha>1$ between $\mathcal{N}(a,\sigma^{2})$ and $\mathcal{N}(0,\sigma^{2})$ is
Therefore, the 1-D Gaussian shift envelope is
Substituting $z=\tilde{\Delta}^{u}_{\infty}(\theta,s_{i},s_{j})$ into (30) yields
$\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}\mathrm{Q}\bigr)\;\leq\;\frac{\alpha}{2\sigma^{2}}\,\bigl(\tilde{\Delta}_{\infty}^{u}(\theta,s_{i},s_{j})\bigr)^{2}.$ (31)
By Definition 4.1, the $\omega$-average sliced Rényi divergence is
$\mathtt{ASD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})=\int_{\mathbb{S}^{d-1}}\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}\mathrm{Q}\bigr)\,d\omega(u).$
Using the bound (31) and the linearity of the integral, we have
$\mathtt{ASD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})$ $\leq\int_{\mathbb{S}^{d-1}}\frac{\alpha}{2\sigma^{2}}\,\bigl(\Delta^{\infty}_{u}(\theta,s_{i},s_{j})\bigr)^{2}\,d\omega(u)$
$=\frac{\alpha}{2\sigma^{2}}\int_{\mathcal{U}}\bigl(\Delta^{\infty}_{u}(\theta,s_{i},s_{j})\bigr)^{2}\,d\omega(u)$
$=\frac{\alpha}{2\sigma^{2}}\,\Delta^{2}(\theta,s_{i},s_{j}).$
Step 3. By definition of $\overline{\Delta}^{2}$,
$\Delta^{2}(\theta,s_{i},s_{j})\leq\overline{\Delta}^{2}$
for all $\theta,(s_{i},s_{j})$. Therefore
@@
$\mathtt{ASD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})\leq\frac{\alpha}{2\sigma^{2}}\,\overline{\Delta}^{2}.$
With the choice
@@
$\sigma^{2}=\frac{\alpha\,\overline{\Delta}^{2}}{2\epsilon},$
we obtain, for every $\theta,(s_{i},s_{j})$,
$\mathtt{ASD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{s_{i}}\,\big\|\,\mathcal{M}^{\theta}_{s_{j}}\bigr)\leq\epsilon.$
Since $\theta$ and $(s_{i},s_{j})$ were arbitrary, this is exactly the $(\alpha,\epsilon,\omega)$-Ave-SRPE guarantee in $(\mathcal{S},\mathcal{Q},\Theta)$. ∎
Appendix H Proof of Theorem 4.8
Fix $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$, and fix a slice profile $\{\mathcal{U},\omega\}$. Let $F^{\theta}_{s}$ denote the pre-noise distribution of $f(X)$ given $(S=s,\theta)$, and let
$\mathrm{P}:=\mathcal{M}^{\theta}_{s_{i}}=F^{\theta}_{s_{i}}*\zeta,\quad\textup{and}\quad\mathrm{Q}:=\mathcal{M}^{\theta}_{s_{j}}=F^{\theta}_{s_{j}}*\zeta.$
@@
where $z^{u}_{\infty}(\theta,s_{i},s_{j})$ upper-bounds the 1-D shift (in $W_{\infty}$) between the pre-noise probability measures $\Psi^{u}_{\#}F^{\theta}_{s_{i}}$ and $\Psi^{u}_{\#}F^{\theta}_{s_{j}}$. By definition,
$z^{u}_{\infty}(\theta,s_{i},s_{j})\leq\tilde{\Delta}^{u}_{\infty}(\theta,s_{i},s_{j}):=W_{\infty}\bigl(\Psi^{u}_{\#}F^{\theta}_{s_{i}}\,\big\|\,\Psi^{u}_{\#}F^{\theta}_{s_{j}}\bigr),$
@@
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}):=\frac{1}{\alpha-1}\log\int_{\mathbb{S}^{d-1}}\exp\Bigl((\alpha-1)\,\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}\mathrm{Q}\bigr)\Bigr)d\omega(u).$
@@
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})$ $=\frac{1}{\alpha-1}\log\int_{\mathbb{S}^{d-1}}\exp\Bigl((\alpha-1)\,\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}\mathrm{Q}\bigr)\Bigr)d\omega(u)$

### v3

@@ Appendix H Minimality and Attainability of HUC
we obtain
$\mathtt{D}_{\alpha}(\widetilde{\zeta}_{z}\|\widetilde{\zeta}_{z^{\prime}})\leq R_{\alpha}(\widetilde{\zeta},|z-z^{\prime}|).$
Combining with (53) yields, for $\pi_{\varepsilon}$-a.e. $(z,z^{\prime})$,
$\mathtt{D}_{\alpha}(\widetilde{\zeta}_{z}\|\widetilde{\zeta}_{z^{\prime}})\leq R_{\alpha}(\widetilde{\zeta},z_{\infty}^{u}+\varepsilon).$
Step 3 (From pointwise to mixtures via DPI). Define probability measures on $\mathbb{R}^{2}\times\mathbb{R}$ by
$\overline{P}_{\varepsilon}(dz,dz^{\prime},dy)=\pi_{\varepsilon}(dz,dz^{\prime})\,\widetilde{\zeta}_{z}(dy),$
$\overline{Q}_{\varepsilon}(dz,dz^{\prime},dy)=\pi_{\varepsilon}(dz,dz^{\prime})\,\widetilde{\zeta}_{z^{\prime}}(dy).$
Their $y$-marginals are exactly the desired mixtures:
$(\overline{P}_{\varepsilon})_{Y}=\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}},\quad(\overline{Q}_{\varepsilon})_{Y}=\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}.$
Since marginalization is a data-independent post-processing, DPI implies
@@
$\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}\ \big\|\ \Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}\right)\leq\mathtt{D}_{\alpha}(\overline{P}_{\varepsilon}\|\overline{Q}_{\varepsilon}).$
Moreover, conditioning on $(z,z^{\prime})$ and using the standard expression for Rényi divergence of a mixture with a common mixing law,
@@
$\mathtt{D}_{\alpha}(\overline{P}_{\varepsilon}\|\overline{Q}_{\varepsilon})=\frac{1}{\alpha-1}\log\int\exp\!\Big((\alpha-1)\mathtt{D}_{\alpha}(\widetilde{\zeta}_{z}\|\widetilde{\zeta}_{z^{\prime}})\Big)\,\pi_{\varepsilon}(dz,dz^{\prime}).$
Using the pointwise bound from Step 2 and the monotonicity of $r\mapsto R_{\alpha}(\widetilde{\zeta},r)$ gives
$\mathtt{D}_{\alpha}(\overline{P}_{\varepsilon}\|\overline{Q}_{\varepsilon})\leq R_{\alpha}(\widetilde{\zeta},z_{\infty}^{u}+\varepsilon).$
Therefore,
@@
$\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}\ \big\|\ \Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}\right)\leq R_{\alpha}(\widetilde{\zeta},z_{\infty}^{u}+\varepsilon).$
Finally, letting $\varepsilon\downarrow 0$ yields
@@
$\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{i}}\ \big\|\ \Psi^{u}_{\#}\mathcal{M}^{\theta}_{s_{j}}\right)\leq R_{\alpha}(\widetilde{\zeta},z_{\infty}^{u}).$
Renaming $\widetilde{\zeta}=\Psi^{u}_{\#}\zeta$ (as in the lemma statement) proves the claim. ∎
Appendix L Proof of Theorem 4.5
Fix $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$. Let $\zeta=\mathcal{N}(0,\sigma^{2}I_{d})$ and define
$\mathrm{P}:=\mathcal{M}^{\theta}_{s_{i}}=P_{\theta}^{f,s_{i}}*\zeta,\quad\mathrm{Q}:=\mathcal{M}^{\theta}_{s_{j}}=P_{\theta}^{f,s_{j}}*\zeta.$
Step 1 (per-slice envelope bound). Fix any direction $u\in\mathcal{U}$. Applying Lemma 4.1 to the 1-D projection $\Psi^{u}(y)=\langle y,u\rangle$ yields
@@
$\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathrm{P}\ \big\|\ \Psi^{u}_{\#}\mathrm{Q}\right)\leq R_{\alpha}\!\left(\zeta,\ z^{u}_{\infty}(\theta,s_{i},s_{j})\right),$
where
$z^{u}_{\infty}(\theta,s_{i},s_{j}):=W_{\infty}\!\Bigl(\Psi^{u}_{\#}P_{\theta}^{f,s_{i}},\ \Psi^{u}_{\#}P_{\theta}^{f,s_{j}}\Bigr).$
By the definition of the per-slice $\infty$-Wasserstein sensitivity $\Delta_{\infty}^{u}$ in (6), for every $u\in\mathcal{U}$ we have
$z^{u}_{\infty}(\theta,s_{i},s_{j})\leq\Delta_{\infty}^{u},$
and since $R_{\alpha}(\zeta,\cdot)$ is non-decreasing,
$\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathrm{P}\ \big\|\ \Psi^{u}_{\#}\mathrm{Q}\right)\leq R_{\alpha}\!\left(\zeta,\ \Delta_{\infty}^{u}\right).$ (54)
@@
Step 2 (Gaussian envelope). For $\zeta=\mathcal{N}(0,\sigma^{2}I_{d})$, the projected noise $\Psi^{u}(N)$ is $\mathcal{N}(0,\sigma^{2})$. For any shift $a\in\mathbb{R}$,
$\mathtt{D}_{\alpha}\!\left(\mathcal{N}(a,\sigma^{2})\ \big\|\ \mathcal{N}(0,\sigma^{2})\right)=\frac{\alpha a^{2}}{2\sigma^{2}}.$
Hence the 1-D Gaussian shift envelope satisfies
$R_{\alpha}(\zeta,z)=\sup_{|a|\leq z}\mathtt{D}_{\alpha}\!\left(\mathcal{N}(a,\sigma^{2})\ \big\|\ \mathcal{N}(0,\sigma^{2})\right)=\frac{\alpha}{2\sigma^{2}}z^{2}.$
Substituting this into (54) gives, for all $u\in\mathcal{U}$,
@@
$\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathrm{P}\ \big\|\ \Psi^{u}_{\#}\mathrm{Q}\right)\leq\frac{\alpha}{2\sigma^{2}}\,(\Delta_{\infty}^{u})^{2}.$
Step 3 (average over $\omega$ and choose $\sigma^{2}$). By Definition 4.1,
$\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})=\int_{\mathbb{S}^{d-1}}\mathtt{D}_{\alpha}\!\left(\Psi^{u}_{\#}\mathrm{P}\ \big\|\ \Psi^{u}_{\#}\mathrm{Q}\right)\,d\omega(u).$
Therefore,
@@
$\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})\leq\int_{\mathbb{S}^{d-1}}\frac{\alpha}{2\sigma^{2}}(\Delta_{\infty}^{u})^{2}\,d\omega(u)=\frac{\alpha}{2\sigma^{2}}\,\overline{\Delta}^{2}.$
Choosing $\sigma^{2}=\frac{\alpha\,\overline{\Delta}^{2}}{2\epsilon}$ yields $\mathtt{AveSD}^{\omega}_{\alpha}(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}})\leq\epsilon$ for all $\theta$ and $(s_{i},s_{j})$, i.e., $\mathcal{M}$ satisfies $(\alpha,\epsilon,\omega)$-Ave-SRPP. ∎
Appendix M Proof of Theorem 4.6
Fix $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$, and fix a slice profile $\{\mathcal{U},\omega\}$. Let
$\mathrm{P}:=\mathcal{M}^{\theta}_{s_{i}}=P_{\theta}^{f,s_{i}}*\zeta,\quad\textup{and}\quad\mathrm{Q}:=\mathcal{M}^{\theta}_{s_{j}}=P_{\theta}^{f,s_{j}}*\zeta.$
@@
where $z^{u}_{\infty}(\theta,s_{i},s_{j})$ upper-bounds the 1-D shift (in $W_{\infty}$) between the pre-noise probability measures $\Psi^{u}_{\#}P_{\theta}^{f,s_{i}}$ and $\Psi^{u}_{\#}P_{\theta}^{f,s_{j}}$. By definition,
$z^{u}_{\infty}(\theta,s_{i},s_{j})\leq\tilde{\Delta}^{u}_{\infty}(\theta,s_{i},s_{j}):=W_{\infty}\bigl(\Psi^{u}_{\#}P_{\theta}^{f,s_{i}}\,\big\|\,\Psi^{u}_{\#}P_{\theta}^{f,s_{j}}\bigr),$
@@
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})$
$=\frac{1}{\alpha-1}\log\int_{\mathbb{S}^{d-1}}\exp\Bigl((\alpha-1)\,\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}\mathrm{Q}\bigr)\Bigr)d\omega(u).$
@@
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})$
$=\frac{1}{\alpha-1}\log\int_{\mathbb{S}^{d-1}}\exp\Bigl((\alpha-1)\,\mathtt{D}_{\alpha}\bigl(\Psi^{u}_{\#}\mathrm{P}\,\big\|\,\Psi^{u}_{\#}\mathrm{Q}\bigr)\Bigr)d\omega(u)$